### v1

# Jing Liu EnQi Lian

Abstract Sound Fusion with Unconditioned Inversion Model
Jing Liu EnQi Lian
Abstract
An abstract sound is defined as a sound that does not disclose identifiable real-world sound events to a listener. Sound fusion aims to synthesize an original sound and a reference sound to generate a novel sound that exhibits auditory features beyond mere additive superposition of the sound constituents. To achieve this fusion, we employ inversion techniques that preserve essential features of the original sample while enabling controllable synthesis. We propose novel SDE and ODE inversion models based on DPMSolver++ samplers that reverse the sampling process by configuring model outputs as constants, eliminating circular dependencies incurred by noise prediction terms. Our inversion approach requires no prompt conditioning while maintaining flexible guidance during sampling.
Inversion models, sound fusion, diffusion models, music generation
1 Introduction
Sound fusion seeks to combine an original sound with a reference sound, creating a new auditory output whose acoustic characteristics transcend simple additive mixing of the input components. In this study, sounds are classified into two primary categories: concrete sounds and abstract sounds. Concrete sounds are those that listeners can clearly associate with specific objects, scenes, events, or musical patterns that are intuitively recognizable due to their semantic content. Examples such as running water, crying children, arriving trains, metallic percussion, and burning firewood can be readily identified as concrete. Conversely, abstract sounds are those that do not reveal identifiable real-world sound events or, in musical contexts—particularly in contemporary, electronic, and experimental music clips—do not intuitively disclose recognizable patterns. In essence, the presence of describable, semantically meaningful patterns determines whether a sound is classified as concrete or abstract.
The classification of sounds as abstract or concrete is relative and contingent upon individual cognition and expertise. For instance, the calls of rare insects or distant whale songs may be unfamiliar to the general population and thus perceived as abstract; however, these same sounds are concrete and meaningful to entomologists and marine biologists in their respective fields of expertise.
This research specifically focuses on the fusion of abstract sounds. While concrete sounds offer limited potential for meaningful fusion due to their inherent semantic specificity, fused abstract sounds demonstrate musical potential that serve as material for composition.
While we use ”sound” to distinguish our focus from music, this research technically operates on digital audio as the carrier medium.
The main contributions of this work are as follows:
  • We propose abstract sound fusion as a novel application in creative AI.
  • We introduce a sound fusion approach that synthesizes auditory features from an original sound and a reference sound.
  • We propose both Stochastic Differential Equation (SDE) and Ordinary Differential Equation (ODE) inversion models based on DPMSolver++ samplers that enable this sound fusion capability.
2 Related Works
Inversion is a technique that captures features of real samples in referenced generation, and it is usually employed in real-sample editing, a specific subcategory of generative tasks. Unlike conventional generation processes that create content ex nihilo, editing involves modifying specific attributes and characteristics of existing samples during the sampling phase. When provided with a reference, generation maintains closer adherence to users’ intentions.
Audio editing has evolved from earlier approaches that primarily focused on applications such as voice conversion, singing voice conversion, and timbre transfer, with research concentrated in speech and music domains. The emergence of Text-To-Audio (TTA) systems in recent years has
$\boldsymbol{x}_{t-1}=\sqrt{\bar{\alpha}_{t-1}}\left(\frac{\boldsymbol{x}_{t}-\sqrt{1-\bar{\alpha}_{t}}\cdot\epsilon^{(t)}_{\theta}}{\sqrt{\bar{\alpha}_{t}}}\right)+\sqrt{1-\bar{\alpha}_{t-1}-\sigma_{t}^{2}}\cdot\epsilon^{(t)}_{\theta}+\sigma_{t}\boldsymbol{\epsilon}_{t}$ (1)
broadened the scope of real-sample editing and enhanced precision through prompt manipulation.
These TTA systems typically consist of integrated pipelines comprising pre-trained models, offering diverse entry points for addressing editing challenges. Models such as CLAP (Wu* et al., 2023) and T5 (Raffel et al., 2023), trained to predict samples from text prompts, form the foundation of TTA systems and enable prompt-based editing capabilities.
Recent research in real audio editing has drawn inspiration from image editing approaches (Mokady et al., 2022; Huberman-Spiegelglas et al., 2024; Roich et al., 2022), introducing new editing tasks such as substituting sound events within audio, as well as adding, removing, replacing, and refining concrete audio content (Wang et al., 2023; Xu et al., 2024). Furthermore, the research scope has expanded from music and speech to encompass general audio.
Most pre-existing inversion studies have aimed to reverse samples back to their initial latent codes $x_{T}$, from which sampling begins. This technique originated during the GAN era (Zhu et al., 2018; Xia et al., 2022; Meng et al., 2022) and facilitates various applications, including in-painting and super-resolution. From an editing perspective, inversion addresses the challenges of unexpected changes and unwanted feature preservation that occur when employing sample-to-sample approaches. Research in this area has predominantly focused on computer vision applications, with the audio field subsequently adapting these techniques, though for both domains, inversion research on diffusion models remains relatively limited.
Prior inversion approaches have worked with DDPM (Ho et al., 2020) and DDIM (Song et al., 2022), either approximating inversion through diffusion itself or by reversing the sampling process. (Huberman-Spiegelglas et al., 2024) computed latent codes at subsequent timesteps directly from $x_{0}$, and recent work (Manor & Michaeli, 2024) extended this approach to the audio domain. (Mokady et al., 2022) worked with the DDIM sampling ODE and swapped timesteps for inversion approximation. Within the context of Classifier-Free Guidance (CFG) (Ho & Salimans, 2022), the authors optimized the negative prompt during inversion while preserving the flexibility of the positive prompt. This approach fine-tuned interpolations against diffusion latents as pivots (Roich et al., 2022). However, fine-tuning embeddings for each edit incurs relatively high computational costs. PPAE (Xu et al., 2024) subsequently utilized this approach for audio editing.
We choose the open-source TTA system Stable Audio Open 1.0 (Evans et al., 2024b) to achieve the best possible perceptual quality. Since this system uses DPMSolver++ (Lu et al., 2023) samplers for which no inversion methods have been proposed, we introduce DPMSolver++ ODE and SDE inversion models to fill this gap. Our approach preserves prompt flexibility, including negative prompts during sampling. Negative prompts, previously termed unconditioned prompts, were originally used as internal mechanisms to introduce variety, but now provide opposing guidance directions in many modern generation models (Rombach et al., 2021; Evans et al., 2024a). We evaluate our inversion latents regarding identity preservation and prompt responsiveness, and empirical evidence demonstrates that latents computed by our models achieve a balance between both objectives.
We first review the sampling process of a diffusion model in 1. It yields a sample at the previous timestep $x_{t-1}$ from $x_{t}$, where $\sigma_{t}=\eta\sqrt{\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}}\sqrt{1-\frac{\bar{\alpha}_{t}}{\bar{\alpha}_{t-1}}}$. Setting $\eta=0$ results in a deterministic DDIM scheme. Setting $\eta\in(0,1]$ results in a stochastic DDIM scheme that interpolates between deterministic sampling and DDPM. When $\eta=1$, the stochastic DDIM is equivalent to DDPM.
2.1 DDIM Inversion
(Mokady et al., 2022) proposed an inversion model based on deterministic DDIM. The authors omitted the stochastic term from Eq. 1 and reversed the timestep direction for inversion. The latent code of the next timestep is then computed as:
$\boldsymbol{x}_{t}:=\sqrt{\frac{\bar{\alpha}_{t}}{\bar{\alpha}_{t-1}}}\boldsymbol{x}_{t-1}+\left(\sqrt{\frac{1-\bar{\alpha}_{t}}{\bar{\alpha}_{t}}}-\sqrt{\frac{1-\bar{\alpha}_{t-1}}{\bar{\alpha}_{t-1}}}\right)\epsilon^{(t-1)}_{\theta}$
They further fine-tuned the inverted latents by computing the loss against latents produced from the diffusion process, optimizing negative prompt embeddings initialized to null, $\psi(^{\prime\prime})$.
2.2 DDPM Inversion
Huberman-Spiegelglas et al. (2024) claimed that continuous noise vectors $\boldsymbol{\epsilon}_{t}$ were tightly correlated, which led to unpredictable manipulation for image editing. They found that computing directly from $\boldsymbol{x}_{0}$ during the forward process imprinted the original signal better on latents. They first computed latents at each timestep with a diffusion process, and altered the noise accumulated from timestep $1$ to $t$ to independent Gaussian noise:
$\boldsymbol{x}_{t}=\frac{\sigma_{t}}{\sigma_{s}}e^{-(\lambda_{t}-\lambda_{s})}\boldsymbol{x}_{s}+2\alpha_{t}\int_{\lambda_{s}}^{\lambda_{t}}e^{-2(\lambda_{t}-\lambda)}\boldsymbol{x}_{\theta}^{(\lambda)}\text{d}\lambda+\sqrt{2}\sigma_{t}\int_{\lambda_{s}}^{\lambda_{t}}e^{-(\lambda_{t}-\lambda)}\text{d}\omega_{\lambda}$ (2)
$\boldsymbol{x}_{r}=\frac{2r_{1}(\sigma_{r}^{2}+1)}{\alpha_{t}(1-e^{-2h})}\left(\boldsymbol{x}_{t}-\frac{\sigma_{t}}{\sigma_{s}}e^{-h}\boldsymbol{x}_{s}+\frac{\alpha_{t}(1-2r_{1})(1-e^{-2h})}{2r_{1}}\boldsymbol{x}_{\theta}^{(s)}+\frac{\sigma_{r}\alpha_{t}(1-e^{-2h})}{2r_{1}\sqrt{\sigma_{r}^{2}+1}}\epsilon^{(r)}_{\theta}\right)$ (3)
$\boldsymbol{x}_{t}=\sqrt{\bar{\alpha}_{t}}\boldsymbol{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\boldsymbol{\epsilon}_{t},\boldsymbol{\epsilon}_{t}\sim\mathcal{N}(0,I)$ (4)
Next, they rearranged Eq. 4 to solve for the noise map $\boldsymbol{\epsilon}_{t}\sim(\boldsymbol{x}_{t},\epsilon^{(t)}_{\theta}(\boldsymbol{x}_{t},\psi))$ at each timestep during sampling, given the diffused latents. To reduce accumulated error, the authors then updated $\boldsymbol{x}_{t-1}$ through the DDPM sampling equation.
3 Methods
In this section, we first introduce the notation employed throughout this paper. We then demonstrate that the noise prediction term introduces circular dependency, leading to numerically unstable fine-tuning. Subsequently, we propose inversion models that reconstruct the original audio and fine-tune Stable Audio Open to reconstruct reference audio. Finally, we present the audio fusion algorithm.
3.1 Notation and Definitions
In DDIM and DDPM, $\alpha$ represents the portion of the signal, and $\bar{\alpha}$ is the cumulative product of $\alpha$, where $\bar{\alpha}_{t}=\prod_{s=1}^{t}\alpha_{s}$. The notation for $\alpha$ and $\bar{\alpha}$ varies in the original papers; however, we use them consistently throughout this paper.
In DPMSolver++, regarding the notation of timesteps $t$ and $s$, $t$ falls within the range $[0,s]$ for $s>0$. Specifically, $t$ precedes $s$ temporally. Lu et al. (2022) transformed the diffusion ODE from the time domain to the half log-SNR domain, which is represented by $\lambda$ and defined as $\lambda:=\log(\frac{\alpha_{t}}{\sigma_{t}})$.
The relation between $\alpha$ and $\sigma$ is given by $\alpha^{2}_{t}+\sigma^{2}_{t}=1$. The model output is defined as follows:
$x^{(t)}_{\theta}=\frac{\boldsymbol{x}_{t}-\sigma_{t}\cdot\epsilon^{(t)}_{\theta}}{\alpha_{t}}$ (5)
$\epsilon^{(t)}_{\theta}$ is an abbreviation for $\epsilon^{(t)}_{\theta}(\boldsymbol{x}_{t},\psi)$, which denotes a network that takes the latent and prompt embeddings $\psi$ as input conditions and outputs the predicted noise at timestep $t$.
$x^{(t)}_{\theta}$ is an abbreviation for $x^{(t)}_{\theta}(\boldsymbol{x}_{t},\epsilon^{(t)}_{\theta}(\boldsymbol{\boldsymbol{x}_{t}},\psi))$. It takes the latent code and predicted noise as inputs at timestep $t$.
3.2 Predicted Noise Introduces Circular Dependency
Lu et al. (2023) proved that many high-order samplers are unstable at large guided scales, which amplified the derivatives of the noise prediction model and affect the convergence range of the ODE solver. The authors proposed replacing the noise prediction term with the model output $x_{\theta}$ to achieve more stable convergence.
Taking a general SDE sampler as an example, given $s>0$, the latent at the previous timestep $t\in[0,s]$ is denoted as shown in Eq. 2.
Previous research performed inversion on predicted noise. However, expanding the model output with the predicted noise expression results in circular dependency. To address this circular dependency issue intuitively, we substitute the general formula in Eq. 2 into the second-order SDE case and target the latent at the next timestep $r$, as shown in Eq. 3.
From this example, $\epsilon_{\theta}$ takes $\boldsymbol{x}_{r}$ as its input, which is the variable we are solving for on the left side.
In this example, $\epsilon_{\theta}$ takes $\boldsymbol{x}_{r}$ as its input, which is the variable we are solving for on the left side. This circular dependency occurs in first- and higher-order methods for both SDE and ODE samplers. Fine-tuning on latent $\boldsymbol{x}_{r}$ becomes difficult because the outer scaling factor grows extremely large during the latter stages of inversion when the maximum $\sigma$ is set to its default value. Consequently, small error accumulation leads to numerically unstable training.
3.3 Model Output as Constant
DPMSolver++ introduced model output to replace the predicted noise term, which provides us with a sampling perspective: for every step to the previous timestep, we desire $\epsilon_{\theta}$ to predict a noise map that perfectly matches the latent such that, given both conditions, the model output is exactly the same as our target sample for inversion. Hence, by assigning the encoded $\boldsymbol{x}_{0}$ of the real sample to $x_{\theta}$ in Eq. 2 and removing the stochastic term, the latent at the next timestep $\boldsymbol{x}_{t}$ is given by:
$\boldsymbol{x}_{t}=\frac{\sigma_{t}}{\sigma_{t-1}}e^{\lambda_{t-1}-\lambda_{t}}\left(\boldsymbol{x}_{t-1}-\alpha_{t-1}(1-e^{-2(\lambda_{t-1}-\lambda_{t})})\boldsymbol{x}_{0}\right)$ (6)
Higher-order integral terms are thus eliminated from the formula, where only the schedulers are considered. Apart from the stochastic term, the drift terms of the ODE variation also differ from those of the SDE. The general ODE formula is as follows:
$\boldsymbol{x}_{t}=\frac{\sigma_{t}}{\sigma_{s}}\boldsymbol{x}_{s}+\sigma_{t}\int_{\lambda_{s}}^{\lambda_{t}}e^{\lambda}x_{\theta}^{(\lambda)}\text{d}\lambda$ (7)
Assigning $\boldsymbol{x}_{0}$ to $x_{\theta}$, the integral is simplified as in the SDE formula. The latent code $\boldsymbol{x}_{t}$ yields:
$\boldsymbol{x}_{t}=\frac{\sigma_{t}}{\sigma_{t-1}}\left(\boldsymbol{x}_{t-1}-\sigma_{t-1}(e^{\lambda_{t-1}}-e^{\lambda_{t}})\boldsymbol{x}_{0}\right)$ (8)
Here, we have inversion models for both SDE and ODE variants of DPMSolver++. We can further derive the noise map at $t$ from Eq. 5:
$\boldsymbol{\epsilon}^{(t)}=\frac{\boldsymbol{x}_{t}-\alpha_{t}\boldsymbol{x}_{0}}{\sigma_{t}}$ (9)
3.4 Audio Fusion Algorithm
Through inversion, we obtain latents that preserve features of the original audio at different noise levels. To fuse with another audio source—namely, the reference audio—we need to capture its features as well. Therefore, we fine-tune the model to memorize the reference audio along with its corresponding prompt. We do not require generalization from this fine-tuning process; rather, our objective is to reconstruct the reference audio faithfully from the prompt.
The fusion process is a modified sampling process. The algorithm operates in two phases: in the first half of fusion, the reference audio prompt dominates to produce interpolations with implicit features of the reference audio. In the second half, we inject the inverted latent into the sampler.
Each term holds explicit significance at the moment of injection. As shown in Eq. 2, the first term is where the inverted interpolation is injected, while the second term holds the interpolations guided by the reference prompt. Collectively, these two terms represent an addition of two audio sources, but the third term provides flexibility for fusion.
The algorithm uses low-order sampling at the beginning of the sampling process to warm up for higher-order sampling, and employs first-order sampling near the end of the sampling process to reduce the impact of numerical instability. Algorithm 1 demonstrate second-order sampling as an example. We ensure that the first- and second-order inversion and sampling remain consistent across timesteps. Additionally, we incorporate the CFG mechanism into the editing process. Let $\boldsymbol{x}$ denote the latent at the current timestep and $\boldsymbol{x}_{i-1}$ denote the latent at the previous timestep.
Algorithm 1 Audio Fusion
  Input: prompt of the reference audio ${}^{\prime}C^{\prime}$, negative prompt ${}^{\prime}U^{\prime}$, inverted latent space $\{\hat{x}\}_{\text{org}}^{i=1,...T}$, injection timestep $t_{j}$, CFG guidance scale $\omega$, warm-up timestep $n_{s}$, cool-down timestep $n_{f}$;
  Output: fused audio $\mathcal{B}$.
  Known: noise prediction network $\epsilon_{\theta}$, noise $z$.
  Project ${}^{\prime}C^{\prime}$ to the embedding $\psi(^{\prime}C^{\prime})$;
  Project ${}^{\prime}U^{\prime}$ to the embedding $\psi(^{\prime}U^{\prime})$;
  Initialize latent $x$.
  for $i=T,...,1$ do
     if $i==t_{j}$ then
        Assign $x\leftarrow\hat{x}_{i}$
     else
        Assign $x\leftarrow x_{i}$
     end if
     Compute the noise residue $\epsilon_{\theta}(x,\psi(^{\prime}C^{\prime}))$;
     Compute the noise residue $\epsilon_{\theta}(x,\psi(^{\prime}U^{\prime}))$;
     Compute the CFG noise residue $\epsilon_{\theta}(x)=\omega\epsilon_{\theta}(x,\psi(^{\prime}C^{\prime}))+(1-\omega)\epsilon_{\theta}(x,\psi(^{\prime}U^{\prime}))$;
     Compute $x_{\theta}(x,\epsilon_{\theta}(x))$ with Eq. 5.
     if $i\geq n_{s}$ or $i<n_{f}$ then
        Compute the first order $x_{i-1}$ using Eq. 2;
     else
        Compute the second order $x_{i-1}$ using Eq. 2.
     end if
     Assign $x_{i}\leftarrow x$.
  end for
  Decode $x_{i-1}$ with the VAE decoder to obtain the fused audio $\mathcal{B}$.
  return $\mathcal{B}$
4 Experiments and Results
The potential for an inverted latent to reconstruct the original sample must be measured perceptually. Metrics evaluating distance or similarity between vectors, including MSE and cosine similarity, do not necessarily correlate with auditory perceptual similarity, which is our primary concern. While perceptual measurements exist for speech, these metrics are tailored to speech-specific characteristics and do not translate well to general audio. Significant differences exist between speech and general audio regarding frequency bands, harmonic content, and spectral characteristics. Most notably, speech typically occupies a narrower frequency range, whereas general audio spans the full audible spectrum and includes diverse timbral qualities, complex textures, and non-harmonic elements that speech-based perceptual metrics cannot adequately capture. To the best of our knowledge, established metrics for measuring perceptual similarity in general audio are lacking. Therefore, we employ spectrograms to provide intuitive visual analysis in lieu of numerical assessment. [^1] All spectrograms presented in this paper are generated using Sonic Visualiser software. The corresponding audio is available online. [^2] https://github.com/Jng-L/SoundFusion
(a) Employ inverted latents from our method.
(b) Employ latents with noise added from forward diffusion.
(c) Employ noisy latents with stochastic terms preserved from our method.
Figure 2: Second-order SDE reconstruction with null prompt guidance. Columns 1-3: intervention at t=0, 80, 140. Column 4: original audio through VAE.
4.1 Original Audio Reconstruction
To edit a real sample, text prompts serve as critical conditions to guide the sampling process, requiring a balance between identity preservation in latents and responsiveness to prompt guidance. Before fusion, we first examine the resistance of latent codes produced by our inversion models to prompt conditioning. We use null prompt embeddings in all experiments in this subsection.
4.1.1 Necessity of Omitting Stochastic Terms
Stochastic terms are added back during sampling because they help preserve the potential for editing. Let $\tilde{x}_{t}$ be given by Eq. 6; then the comparisons are as follows:
  1. (a)
    $x_{t}=\tilde{x}_{t}$
  2. (b)
    $x_{t}=x_{0}+\sigma_{t}\epsilon$
  3. (c)
    $x_{t}=\tilde{x}_{t}-\sigma_{t}\sqrt{e^{2(\lambda_{t-1}-\lambda_{t})}-1}\epsilon_{t}$
The first observation is illustrated in Figure 2: any form of noise added during inversion, including retaining stochastic terms or adding noise as in forward diffusion, results in a very noisy sampling phase. This observation demonstrates that omitting noise in inversion is reasonable, and that approximating inversion with diffusion produces samples of unacceptable quality with DPMSolver++. The original audio is reconstructed through VAE encoding and decoding to eliminate the influence of the VAE itself.
4.1.2 Potential of Initial Latent
Figure 3: Comparison of audio generated from initial latents under different methods. Prompt: ”An airplane takes off.” (a) Original audio through VAE. (b) Sample from scratch. (c) Our inversion approach. (d) Audio-to-audio adaptation on Stable Audio Open 1.0. (e) Built-in audio-to-audio of Stable Audio 2.0.
We evaluate the SDE inversion model independently on the initial latent. This evaluation is based on the assumption that, apart from the initial latent representation, the model, sampling process, parameters, and other conditions remain the same. The accuracy with which the initial latent representation captures the characteristics of the original audio determines whether the interpolations in the generation process can precisely produce sound containing the original audio’s features and the degree of fidelity. We compare our method with existing approaches. The comparison methods include:
  • Sampling from scratch.
  • Image-to-image methods from the visual domain, adapted to the Stable Audio Open workflow for audio-to-audio editing.
  • The audio-to-audio capabilities of Stable Audio 2.0.
In the open-source audio-to-audio method, we encode the original audio through a VAE and incorporate it into randomly initialized latent representations. To ensure fair comparison, we maintain consistency across all methods using Stable Audio Open, including audio-to-audio, generating from scratch, and our inversion model. Since our objective is to examine the potential of initial latent representations, we deliberately avoid introducing additional adaptive editing schemes during the sampling process. Specifically, we do not intervene in the inverted latent representation throughout sampling but maintain methodological consistency by solely modifying the initial latent assignment. For evaluation, we employ common knowledge prompts, those with sufficient training data during pre-training, as text inputs. We apply identical text prompts and original audio (except for generating from scratch) across all comparative methods, ensuring that the same sampler and parameters are applied to maintain experimental validity.
Stable Audio 2 is a commercial version, and both its built-in audio-to-audio function and the sampling process remain black boxes to us. While it claims to have incorporated this functionality on top of the previous version (Evans et al., 2024a), it is unclear whether it employs the same methodology as the open-source approach. Additionally, this version demonstrates improved model capabilities compared to the previous version, so we include it as an additional reference point.
There are subtle differences among the generated audio from the open-source model, and it is practically impossible to distinguish the original sound components across all five samples (see Figure 3). As expected, audio generated by other methods should exhibit greater similarity to the original sound than samples generated from scratch; however, all outputs of Stable Audio Open sound similar regardless of the original audio content. Though the sounds generated by the open-source and commercial versions vary, both adhere to the text prompt ”An airplane takes off.”
As discussed in the methodology section, this phenomenon arises due to the noise control coefficient and signal strength coefficient, which make the noise guided by text prompts dominate generation during the early sampling stages. This is particularly evident in generation from scratch, where the influence of the initial latent representation is relatively weak. These findings indicate that additional controls over the sampling trajectory need to be integrated into the sampling process to achieve ideal reconstruction results.
4.1.3 Latent Intervention Across Sampler Orders
Figure A.2 demonstrates reconstructed samples when intervening with inverted latent codes at timesteps 0, 80, 120, 140, and 180, compared to the sample $x_{0}$ in the last column. The three rows employ first-, second-, and third-order samplers respectively. First-order sampling is equivalent to DDIM. It is evident that features illustrated in the third-order reconstruction are relatively vague compared to the other two. Although we insert the latent at certain timesteps, higher-order methods compute with interpolations, therefore, the influence of these interpolations persists. Consequently, the intervened latents do not dominate the reconstruction to the same extent as in first- and second-order sampling. At the end of sampling, sigma, the noise scheduler, becomes very small, and the predicted noise cannot compete with the inserted latents that preserve most of the features. This indicates that initially, the prompt controls the generation process, where sigma is fairly large.
4.2 Reference Audio Reconstruction
The audio reconstructed through model fine-tuning restores the reference sound fairly completely, but some loss and confusion of details still exist. As shown in the spectrograms of Figure 4, the reconstruction process generates a layer of blurry background noise, with the sound’s stylistic features enhanced over this background noise. This phenomenon is most evident in the low-frequency portions. As we analyzed earlier, the VAE is designed to prioritize audibility and tends to smooth the regions around sampling points, leading to such artifacts in the sound. While the sound lacks clarity in details, it is acceptable overall. Therefore, we believe that the noise prediction module has acquired the capability to reconstruct text descriptions into audio with reference sound characteristics.
Figure 4: Spectrograms of (a) reference audio and (b) reconstructed reference audio.
4.3 Audio Fusion
We observe that inverted latent representation intervention at timesteps within a certain interval produces audio with optimal fusion results, indicating the existence of a fusion window. Within this window, as the intervention timestep decreases, the fused audio gradually transitions from resembling the reference audio to resembling the original audio in terms of auditory characteristics. Specifically, fusion represents a process of mutual balance between the reference sound and the original sound, where identifying the equilibrium point between the two within the window becomes crucial. In this section, we present the results of audio fusion and subjective evaluation of the fusion effectiveness.
(a) Set 1.
(b) Set 2.
(c) Set 3.
(d) Set 4.
(e) Set 5.
Figure 6: Sound fusion results. The first, second, and third columns represent the original sound, reference sound, and fused sound, respectively.
4.3.1 Results
We used text descriptions as control conditions for the noise prediction network, intervening with inverted latents during the sampling process at regular intervals of 10 timesteps, starting from $t=0$. When intervention occurs near the end of sampling, the sound is almost indistinguishable from the original sound, and the reference sound components are difficult to recognize. We therefore intervened with latent representations in the middle-to-late stages of the sampling process to achieve optimal fusion results. Figure 6 shows fusion examples.
4.3.2 Perceptual Evaluation of Fusion Effects
We received a total of 48 valid questionnaires, with most respondents frequently participating in music activities. Professional and vocational practitioners constitute the largest proportion, as shown in Figure 7.
Figure 7: Distribution of respondents’ level of music engagement.
We observed in our experiments that as the intervention timestep decreased, the auditory similarity of the fused sound gradually transitioned from resembling the reference sound to being more similar to the original sound. We selected fusion examples where the intervention timesteps are positioned within the fusion window, and the similarity distribution from subjective evaluation aligns well with our experimental observations. We used the fusion examples from the previous section to query respondents. In Figure 8, ”Audio 1” represents the original audio, ”Audio 2” represents the reference audio, and ”Audio 3” represents the fused audio; the first through fifth audio sets correspond to Set 1 through Set 5 in Figure 6, respectively. Overall, respondents rated the fusion effects favorably, indicating that we successfully achieved the goal of sound fusion.
(a) Set 1.
(b) Set 2.
(c) Set 3.
(d) Set 4.
(e) Set 5.
Figure 8: Auditory perceptual similarity evaluation results.
5 Conclusion
In this paper, we propose inversion models based on DPMSolver++ with two variations: SDE and ODE. We make the integral term as constant, thereby avoiding the circular dependency of expanding the noise prediction within the sampling formula. Additionally, the deterministic formulation allows for straightforward computation of the noise map at each inversion step.
Our approach retains the flexibility of prompt guidance for both positive and negative guidance. It is computationally efficient. We explain the necessity of removing stochastic terms from the inversion model, demonstrating that approximating inversion with forward diffusion is not applicable for DPMSolver++ sampling phase.
Previous inversion approaches typically computed only the initial latent, and our experiments show that the initial latent alone is insufficient to capture features of the original sample. Therefore, for editing applications, it must be combined with additional control approaches, such as optimized prompt embeddings that control the sampling trajectory, or by intervening in the sampling phase as we propose. In the latter case, we do not necessarily need to invert all the way to the final timestep.
Inversion is employed to reconstruct the auditory features of the original sound, while fine-tuning helps preserve the features of the reference sound through prompting. An optimal interval exists during which intervention achieves effective fusion. As the timestep decreases, the auditory features of the fused sound shift from those of the reference sound to those of the original sound. Higher-order samplers provide a wider fusion interval. Due to the lack of perceptual metrics for general audio, and because our sound fusion approach does not conform to specific physical constraints, as indicated by participant feedback in Appendix B, objective evaluation of our fusion results presents significant challenges. Therefore, we primarily rely on subjective auditory evaluation. The results validate the effectiveness of the sound fusion and inversion methods proposed in this work.
Appendix A Latent Intervention across Sampler Orders: Additional Results
This section provides comprehensive visual results for the latent intervention experiments across different sampler orders discussed in Section 4.1.3. Figure A.2 shows detailed spectrograms comparing first-, second-, and third-order sampling with interventions at various timesteps.
(a) Audio reconstructed by first-order sampling.
(b) Audio reconstructed by second-order sampling.
(c) Audio reconstructed by third-order sampling.
Figure A.2: SDE reconstruction after inversion with null prompt guidance. Columns 1-5: intervention at t=0, 80, 120, 140, 160. Column 6: original audio through VAE.
Appendix B Subjective Analysis on Fusion Results
Fusion is achieved through music generation models, making it difficult to control granularity at the physical or acoustic level. Professional electronic music analysis reveals distinct fusion patterns across the examples.
In the first example, the original audio concentrates in high frequencies, while the reference audio shows evenly distributed frequencies with prominent highs. The fused audio resembles the reference audio’s frequency distribution but adopts the ”melodic direction” of the original audio.
In the second example, the original audio exhibits evenly distributed frequencies with prominent highs and minimal reverb; the reference audio has obvious melodic characteristics with concentrated high tones and some reverb; the fused audio has significant reverb and combines melodies from both sources.
For the third example, the original audio has obvious rhythm, mainly percussion-based with repetitive melody; the reference audio has a wide melodic range with almost no rhythm and sustained low frequencies; the fused audio combines the rhythmic nature of the original audio with the sustained and repetitive characteristics of the reference audio’s low frequencies.
In the fourth example, the original audio has a wide melodic range; the reference audio is essentially pure noise; the fused audio basically repeats the original audio’s melody but incorporates substantial low-frequency noise from the reference.
In the fifth example, the original audio is primarily rhythmic with percussion with layered melody; the reference audio is pure noise with spatial characteristics; the fused audio combines the rhythm and melody of the original with the spatial characteristics of the reference, amplifying these effects.
Appendix C Derived Noise Maps
At the end of Section 3.3, we derived the noise map at timestep $t$ given $x_{0}$ and $x_{t}$ in Eq. 9. Below, we demonstrate the relationships between inverted latent interpolations and derived noise maps, as well as the interpolations in the sampling process that utilize the inverted latent and its corresponding noise map at each step:
(a) Interpolations during inversion (time steps increasing from left to right).
(b) Noise maps computed from latent interpolations at the same timestep of the inversion.
(c) Restoration of the original sound through sampling using inverted latent interpolations and noise maps (timesteps decreasing from left to right).
Figure C.2: Sampling from inverted latents and derived noise maps, such that the sampling trajectory coincides with the inversion trajectory.
Appendix D VAE Blurs Acoustic Details
Initially, the audio is encoded by the Variational AutoEncoder (VAE) to derive $x_{0}$. Figure D.1 presents a comparative analysis: the left panel depicts the spectrogram of the original audio signal; the center shows the spectrogram of audio reconstructed through VAE encoding-decoding processes; the right shows audio reconstructed from inverted $x_{t}$ using null prompt guidance. While GANs typically produce high-definition samples, VAEs generate comparatively smoother outputs. Constrained by these limitations, VAE encoding-decoding procedures result in perceptible attenuation of acoustic details, manifested in the spectrogram as indistinct boundaries and homogenized frequency band stratification. This phenomenon derives from a fundamental VAE property: latent space continuity. To enhance generation fluidity and minimize artifacts and distortion, proximate points within the VAE latent space decode to similar outputs, though these points may not correspond to actual encoded inputs. The right-side representation, when compared to the center, exhibits additional loss of features in upper frequency bands that possess low energy. Nevertheless, the VAE preserves the macro-structure, with pronounced periodic characteristics remaining intact. The left-center comparison serves to demonstrate inherent model constraints, isolating these from potential effects attributable to the inversion models themselves. Consequently, evaluations in this paper compare reconstructions through VAE encoding-decoding with original samples, rather than comparing with the original samples directly.
Figure D.1: Spectrogram of the original audio (left), spectrogram of audio reconstructed through VAE encoding-decoding (center), and spectrogram of audio reconstructed by sampling from the inverted latent intervened at step 100/ 200 (right).

### v2

@@
-Jing Liu EnQi Lian
-Abstract Sound Fusion with Unconditioned Inversion Model
-Jing Liu EnQi Lian
+Jing Liu, Enqi Lian, Moyao Deng
+Abstract Sound Fusion with Unconditional Inversion Models
+Jing Liu, Enqi Lian, Moyao Deng
+Department of Music Artificial Intelligence and Music Information Technology
+Central Conservatory of Music
+Beijing, China
+{jing.liu, enqilian, 21A017}@mail.ccom.edu.cn
+Correspondence to Jing Liu.
 Abstract
@@
 Our inversion approach requires no prompt conditioning while maintaining flexible guidance during sampling.
-Inversion models, sound fusion, diffusion models, music generation
 1 Introduction
@@
 Concrete sounds are those that listeners can clearly associate with specific objects, scenes, events, or musical patterns that are intuitively recognizable due to their semantic content.
-Examples such as running water, crying children, arriving trains, metallic percussion, and burning firewood can be readily identified as concrete.
+Examples such as running water, crying children, arriving trains, and metallic percussion can be readily identified as concrete.
 Conversely, abstract sounds are those that do not reveal identifiable real-world sound events or, in musical contexts—particularly in contemporary, electronic, and experimental music clips—do not intuitively disclose recognizable patterns.
@@
 For instance, the calls of rare insects or distant whale songs may be unfamiliar to the general population and thus perceived as abstract; however, these same sounds are concrete and meaningful to entomologists and marine biologists in their respective fields of expertise.
-This research specifically focuses on the fusion of abstract sounds.
-While concrete sounds offer limited potential for meaningful fusion due to their inherent semantic specificity, fused abstract sounds demonstrate musical potential that serve as material for composition.
-While we use ”sound” to distinguish our focus from music, this research technically operates on digital audio as the carrier medium.
+This research specifically focuses on the fusion of abstract sounds; such fused abstract sounds demonstrate musical potential that serve as an emergent material for modern composition.
+We use "sound" to distinguish our focus from music, and "audio" as the carrier medium of sound.
 The main contributions of this work are as follows:
   • We propose abstract sound fusion as a novel application in creative AI.
-  • We introduce a sound fusion approach that synthesizes auditory features from an original sound and a reference sound.
-  • We propose both Stochastic Differential Equation (SDE) and Ordinary Differential Equation (ODE) inversion models based on DPMSolver++ samplers that enable this sound fusion capability.
+  • We introduce an audio fusion approach that synthesizes auditory features from an original audio and a reference audio.
+  • We propose both Stochastic Differential Equation (SDE) and Ordinary Differential Equation (ODE) inversion models based on DPMSolver++ samplers that enable this audio fusion capability.
 2 Related Works
@@
 When provided with a reference, generation maintains closer adherence to users’ intentions.
-Audio editing has evolved from earlier approaches that primarily focused on applications such as voice conversion, singing voice conversion, and timbre transfer, with research concentrated in speech and music domains.
-The emergence of Text-To-Audio (TTA) systems in recent years has
-$\boldsymbol{x}_{t-1}=\sqrt{\bar{\alpha}_{t-1}}\left(\frac{\boldsymbol{x}_{t}-\sqrt{1-\bar{\alpha}_{t}}\cdot\epsilon^{(t)}_{\theta}}{\sqrt{\bar{\alpha}_{t}}}\right)+\sqrt{1-\bar{\alpha}_{t-1}-\sigma_{t}^{2}}\cdot\epsilon^{(t)}_{\theta}+\sigma_{t}\boldsymbol{\epsilon}_{t}$ (1)
-broadened the scope of real-sample editing and enhanced precision through prompt manipulation.
-These TTA systems typically consist of integrated pipelines comprising pre-trained models, offering diverse entry points for addressing editing challenges.
-Models such as CLAP (Wu* et al., 2023) and T5 (Raffel et al., 2023), trained to predict samples from text prompts, form the foundation of TTA systems and enable prompt-based editing capabilities.
-Recent research in real audio editing has drawn inspiration from image editing approaches (Mokady et al., 2022; Huberman-Spiegelglas et al., 2024; Roich et al., 2022), introducing new editing tasks such as substituting sound events within audio, as well as adding, removing, replacing, and refining concrete audio content (Wang et al., 2023; Xu et al., 2024).
-Furthermore, the research scope has expanded from music and speech to encompass general audio.
+Recent research in real audio editing has drawn inspiration from image editing approaches null-text ; edit-friendly ; pivotal_tuning ; SDEdit ; sto_diff_editing , introducing new editing tasks such as substituting sound events within audio, as well as adding, removing, replacing, and refining concrete audio content AUDIT ; PPAE .
+This research scope expands from concrete audio to encompass abstract audio.
 Most pre-existing inversion studies have aimed to reverse samples back to their initial latent codes $x_{T}$, from which sampling begins.
-This technique originated during the GAN era (Zhu et al., 2018; Xia et al., 2022; Meng et al., 2022) and facilitates various applications, including in-painting and super-resolution.
+This technique originated during the GAN era manifold ; GAN_inversion_survey and has since been extended to diffusion models EDICT ; BELM .
 From an editing perspective, inversion addresses the challenges of unexpected changes and unwanted feature preservation that occur when employing sample-to-sample approaches.
-Research in this area has predominantly focused on computer vision applications, with the audio field subsequently adapting these techniques, though for both domains, inversion research on diffusion models remains relatively limited.
-Prior inversion approaches have worked with DDPM (Ho et al., 2020) and DDIM (Song et al., 2022), either approximating inversion through diffusion itself or by reversing the sampling process.
-(Huberman-Spiegelglas et al., 2024) computed latent codes at subsequent timesteps directly from $x_{0}$, and recent work (Manor & Michaeli, 2024) extended this approach to the audio domain.
-(Mokady et al., 2022) worked with the DDIM sampling ODE and swapped timesteps for inversion approximation.
-Within the context of Classifier-Free Guidance (CFG) (Ho & Salimans, 2022), the authors optimized the negative prompt during inversion while preserving the flexibility of the positive prompt.
-This approach fine-tuned interpolations against diffusion latents as pivots (Roich et al., 2022).
+Prior audio inversion approaches have worked with Denoising Diffusion Probabilistic Models (DDPM) DDPM and Denoising Diffusion Implicit Models (DDIM) DDIM , either approximating inversion through diffusion itself or by reversing the sampling process.
+Huberman-Spiegelglas et al.
+edit-friendly computed latent codes at subsequent timesteps directly from $x_{0}$, and recent work zero-shot extended this approach to the audio domain.
+Mokady et al.
+null-text worked with the DDIM sampling ODE and swapped timesteps for inversion approximation.
+Within the Classifier-Free Guidance (CFG) framework CFG , the authors optimized the negative prompt during inversion while preserving the flexibility of the positive prompt.
+This approach fine-tuned interpolations against diffusion latents as pivots pivotal_tuning .
 However, fine-tuning embeddings for each edit incurs relatively high computational costs.
-PPAE (Xu et al., 2024) subsequently utilized this approach for audio editing.
-We choose the open-source TTA system Stable Audio Open 1.0 (Evans et al., 2024b) to achieve the best possible perceptual quality.
-Since this system uses DPMSolver++ (Lu et al., 2023) samplers for which no inversion methods have been proposed, we introduce DPMSolver++ ODE and SDE inversion models to fill this gap.
-Our approach preserves prompt flexibility, including negative prompts during sampling.
-Negative prompts, previously termed unconditioned prompts, were originally used as internal mechanisms to introduce variety, but now provide opposing guidance directions in many modern generation models (Rombach et al., 2021; Evans et al., 2024a).
-We evaluate our inversion latents regarding identity preservation and prompt responsiveness, and empirical evidence demonstrates that latents computed by our models achieve a balance between both objectives.
-We first review the sampling process of a diffusion model in 1.
-It yields a sample at the previous timestep $x_{t-1}$ from $x_{t}$, where $\sigma_{t}=\eta\sqrt{\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}}\sqrt{1-\frac{\bar{\alpha}_{t}}{\bar{\alpha}_{t-1}}}$.
-Setting $\eta=0$ results in a deterministic DDIM scheme.
-Setting $\eta\in(0,1]$ results in a stochastic DDIM scheme that interpolates between deterministic sampling and DDPM.
-When $\eta=1$, the stochastic DDIM is equivalent to DDPM.
-2.1 DDIM Inversion
-(Mokady et al., 2022) proposed an inversion model based on deterministic DDIM.
-The authors omitted the stochastic term from Eq.
-1 and reversed the timestep direction for inversion.
-The latent code of the next timestep is then computed as:
-$\boldsymbol{x}_{t}:=\sqrt{\frac{\bar{\alpha}_{t}}{\bar{\alpha}_{t-1}}}\boldsymbol{x}_{t-1}+\left(\sqrt{\frac{1-\bar{\alpha}_{t}}{\bar{\alpha}_{t}}}-\sqrt{\frac{1-\bar{\alpha}_{t-1}}{\bar{\alpha}_{t-1}}}\right)\epsilon^{(t-1)}_{\theta}$
-They further fine-tuned the inverted latents by computing the loss against latents produced from the diffusion process, optimizing negative prompt embeddings initialized to null, $\psi(^{\prime\prime})$.
-2.2 DDPM Inversion
-Huberman-Spiegelglas et al.
-(2024) claimed that continuous noise vectors $\boldsymbol{\epsilon}_{t}$ were tightly correlated, which led to unpredictable manipulation for image editing.
-They found that computing directly from $\boldsymbol{x}_{0}$ during the forward process imprinted the original signal better on latents.
-They first computed latents at each timestep with a diffusion process, and altered the noise accumulated from timestep $1$ to $t$ to independent Gaussian noise:
-$\boldsymbol{x}_{t}=\frac{\sigma_{t}}{\sigma_{s}}e^{-(\lambda_{t}-\lambda_{s})}\boldsymbol{x}_{s}+2\alpha_{t}\int_{\lambda_{s}}^{\lambda_{t}}e^{-2(\lambda_{t}-\lambda)}\boldsymbol{x}_{\theta}^{(\lambda)}\text{d}\lambda+\sqrt{2}\sigma_{t}\int_{\lambda_{s}}^{\lambda_{t}}e^{-(\lambda_{t}-\lambda)}\text{d}\omega_{\lambda}$ (2)
-$\boldsymbol{x}_{r}=\frac{2r_{1}(\sigma_{r}^{2}+1)}{\alpha_{t}(1-e^{-2h})}\left(\boldsymbol{x}_{t}-\frac{\sigma_{t}}{\sigma_{s}}e^{-h}\boldsymbol{x}_{s}+\frac{\alpha_{t}(1-2r_{1})(1-e^{-2h})}{2r_{1}}\boldsymbol{x}_{\theta}^{(s)}+\frac{\sigma_{r}\alpha_{t}(1-e^{-2h})}{2r_{1}\sqrt{\sigma_{r}^{2}+1}}\epsilon^{(r)}_{\theta}\right)$ (3)
-$\boldsymbol{x}_{t}=\sqrt{\bar{\alpha}_{t}}\boldsymbol{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\boldsymbol{\epsilon}_{t},\boldsymbol{\epsilon}_{t}\sim\mathcal{N}(0,I)$ (4)
-Next, they rearranged Eq.
-4 to solve for the noise map $\boldsymbol{\epsilon}_{t}\sim(\boldsymbol{x}_{t},\epsilon^{(t)}_{\theta}(\boldsymbol{x}_{t},\psi))$ at each timestep during sampling, given the diffused latents.
-To reduce accumulated error, the authors then updated $\boldsymbol{x}_{t-1}$ through the DDPM sampling equation.
+PPAE PPAE subsequently utilized this approach for audio editing.
+We choose the open-source Text-To-Audio (TTA) system Stable Audio Open 1.0 stable_audio_open to achieve the best possible perceptual quality.
+This system uses DPMSolver++ DPMSolver++ samplers, we introduce DPMSolver++ ODE and SDE inversion models in this work.
+Since our approach obviates the need for prompt optimization to constrain the inversion trajectory, it preserves prompt flexibility for future creative applications.
+In this paper specifically, the fusion phase comprises two distinct stages, preceded by fine-tuning the model with the prompt paired with the reference audio.
+During the initial fusion stage, the prompt effectively guides the sampling trajectory toward reconstructing the reference audio; subsequently, following the intervention of the inverted latent of the original audio, we employ a null text prompt to complete the fusion process.
 3 Methods
 In this section, we first introduce the notation employed throughout this paper.
-We then demonstrate that the noise prediction term introduces circular dependency, leading to numerically unstable fine-tuning.
-Subsequently, we propose inversion models that reconstruct the original audio and fine-tune Stable Audio Open to reconstruct reference audio.
+Next, we propose inversion models that reconstruct the original audio and describe the fine-tuning of Stable Audio Open to reconstruct reference audio.
 Finally, we present the audio fusion algorithm.
@@
 In DDIM and DDPM, $\alpha$ represents the portion of the signal, and $\bar{\alpha}$ is the cumulative product of $\alpha$, where $\bar{\alpha}_{t}=\prod_{s=1}^{t}\alpha_{s}$.
-The notation for $\alpha$ and $\bar{\alpha}$ varies in the original papers; however, we use them consistently throughout this paper.
-In DPMSolver++, regarding the notation of timesteps $t$ and $s$, $t$ falls within the range $[0,s]$ for $s>0$.
-Specifically, $t$ precedes $s$ temporally.
-Lu et al.
-(2022) transformed the diffusion ODE from the time domain to the half log-SNR domain, which is represented by $\lambda$ and defined as $\lambda:=\log(\frac{\alpha_{t}}{\sigma_{t}})$.
+Although the notation for $\alpha$ and $\bar{\alpha}$ varies in the original sources, we use them consistently throughout this work.
+[…]
 The relation between $\alpha$ and $\sigma$ is given by $\alpha^{2}_{t}+\sigma^{2}_{t}=1$. The model output is defined as follows:
@@
 It takes the latent code and predicted noise as inputs at timestep $t$.
-3.2 Predicted Noise Introduces Circular Dependency
-Lu et al.
-(2023) proved that many high-order samplers are unstable at large guided scales, which amplified the derivatives of the noise prediction model and affect the convergence range of the ODE solver.
-The authors proposed replacing the noise prediction term with the model output $x_{\theta}$ to achieve more stable convergence.
-Taking a general SDE sampler as an example, given $s>0$, the latent at the previous timestep $t\in[0,s]$ is denoted as shown in Eq. 2.
-Previous research performed inversion on predicted noise.
-However, expanding the model output with the predicted noise expression results in circular dependency.
-To address this circular dependency issue intuitively, we substitute the general formula in Eq.
-2 into the second-order SDE case and target the latent at the next timestep $r$, as shown in Eq.
-3.
-From this example, $\epsilon_{\theta}$ takes $\boldsymbol{x}_{r}$ as its input, which is the variable we are solving for on the left side.
-In this example, $\epsilon_{\theta}$ takes $\boldsymbol{x}_{r}$ as its input, which is the variable we are solving for on the left side.
-This circular dependency occurs in first- and higher-order methods for both SDE and ODE samplers.
-Fine-tuning on latent $\boldsymbol{x}_{r}$ becomes difficult because the outer scaling factor grows extremely large during the latter stages of inversion when the maximum $\sigma$ is set to its default value.
-Consequently, small error accumulation leads to numerically unstable training.
 3.2 Model Output as Constant
-DPMSolver++ introduced model output to replace the predicted noise term, which provides us with a sampling perspective: for every step to the previous timestep, we desire $\epsilon_{\theta}$ to predict a noise map that perfectly matches the latent such that, given both conditions, the model output is exactly the same as our target sample for inversion.
-Hence, by assigning the encoded $\boldsymbol{x}_{0}$ of the real sample to $x_{\theta}$ in Eq.
-2 and removing the stochastic term, the latent at the next timestep $\boldsymbol{x}_{t}$ is given by:
+DPMSolver++ introduced model output to replace the predicted noise term, which provides us with a sampling perspective: at every step toward the previous timestep, we desire the noise prediction network to predict a noise map that perfectly matches the latent such that, given both conditions, the model output exactly equals our target sample for inversion.
+An SDE sampler takes the following form:
+$\boldsymbol{x}_{t}=\frac{\sigma_{t}}{\sigma_{s}}e^{-(\lambda_{t}-\lambda_{s})}\boldsymbol{x}_{s}+2\alpha_{t}\int_{\lambda_{s}}^{\lambda_{t}}e^{-2(\lambda_{t}-\lambda)}x_{\theta}^{(\lambda)}\text{d}\lambda+\sqrt{2}\sigma_{t}\int_{\lambda_{s}}^{\lambda_{t}}e^{-(\lambda_{t}-\lambda)}\text{d}\omega_{\lambda}$ (2)
+Hence, by assigning the encoded $\boldsymbol{x}_{0}$ of the real sample to $x_{\theta}$ in Eq. 2 and removing the stochastic term, the latent at the next timestep $\boldsymbol{x}_{t}$ is given by:
 $\boldsymbol{x}_{t}=\frac{\sigma_{t}}{\sigma_{t-1}}e^{\lambda_{t-1}-\lambda_{t}}\left(\boldsymbol{x}_{t-1}-\alpha_{t-1}(1-e^{-2(\lambda_{t-1}-\lambda_{t})})\boldsymbol{x}_{0}\right)$ (3)
@@
 $\boldsymbol{x}_{t}=\frac{\sigma_{t}}{\sigma_{t-1}}\left(\boldsymbol{x}_{t-1}-\sigma_{t-1}(e^{\lambda_{t-1}}-e^{\lambda_{t}})\boldsymbol{x}_{0}\right)$ (5)
-Here, we have inversion models for both SDE and ODE variants of DPMSolver++. We can further derive the noise map at $t$ from Eq. 5:
-$\boldsymbol{\epsilon}^{(t)}=\frac{\boldsymbol{x}_{t}-\alpha_{t}\boldsymbol{x}_{0}}{\sigma_{t}}$ (9)
 3.3 Audio Fusion Algorithm
@@
 The algorithm operates in two phases: in the first half of fusion, the reference audio prompt dominates to produce interpolations with implicit features of the reference audio.
-In the second half, we inject the inverted latent into the sampler.
-Each term holds explicit significance at the moment of injection.
+In the second half, we intervene the inverted latent into the sampler.
+Each term holds explicit significance at the moment of intervention.
 As shown in Eq.
 2, the first term is where the inverted interpolation is injected, while the second term holds the interpolations guided by the reference prompt.
-Collectively, these two terms represent an addition of two audio sources, but the third term provides flexibility for fusion.
-The algorithm uses low-order sampling at the beginning of the sampling process to warm up for higher-order sampling, and employs first-order sampling near the end of the sampling process to reduce the impact of numerical instability.
-Algorithm 1 demonstrate second-order sampling as an example.
-We ensure that the first- and second-order inversion and sampling remain consistent across timesteps.
-Additionally, we incorporate the CFG mechanism into the editing process.
-Let $\boldsymbol{x}$ denote the latent at the current timestep and $\boldsymbol{x}_{i-1}$ denote the latent at the previous timestep.
-Algorithm 1 Audio Fusion
-  Input: prompt of the reference audio ${}^{\prime}C^{\prime}$, negative prompt ${}^{\prime}U^{\prime}$, inverted latent space $\{\hat{x}\}_{\text{org}}^{i=1,...T}$, injection timestep $t_{j}$, CFG guidance scale $\omega$, warm-up timestep $n_{s}$, cool-down timestep $n_{f}$;
-  Output: fused audio $\mathcal{B}$.
-  Known: noise prediction network $\epsilon_{\theta}$, noise $z$.
-  Project ${}^{\prime}C^{\prime}$ to the embedding $\psi(^{\prime}C^{\prime})$;
-  Project ${}^{\prime}U^{\prime}$ to the embedding $\psi(^{\prime}U^{\prime})$;
-  Initialize latent $x$.
-  for $i=T,...,1$ do
-     if $i==t_{j}$ then
-        Assign $x\leftarrow\hat{x}_{i}$
-     else
-        Assign $x\leftarrow x_{i}$
-     end if
-     Compute the noise residue $\epsilon_{\theta}(x,\psi(^{\prime}C^{\prime}))$;
-     Compute the noise residue $\epsilon_{\theta}(x,\psi(^{\prime}U^{\prime}))$;
-     Compute the CFG noise residue $\epsilon_{\theta}(x)=\omega\epsilon_{\theta}(x,\psi(^{\prime}C^{\prime}))+(1-\omega)\epsilon_{\theta}(x,\psi(^{\prime}U^{\prime}))$;
-     Compute $x_{\theta}(x,\epsilon_{\theta}(x))$ with Eq. 5.
-     if $i\geq n_{s}$ or $i<n_{f}$ then
-        Compute the first order $x_{i-1}$ using Eq. 2;
-     else
-        Compute the second order $x_{i-1}$ using Eq. 2.
-     end if
-     Assign $x_{i}\leftarrow x$.
-  end for
-  Decode $x_{i-1}$ with the VAE decoder to obtain the fused audio $\mathcal{B}$.
-  return $\mathcal{B}$
+Collectively, these two terms represent an addition of two audio sources, while the second and third terms provide flexibility for fusion.
 4 Experiments and Results
-The potential for an inverted latent to reconstruct the original sample must be measured perceptually.
-Metrics evaluating distance or similarity between vectors, including MSE and cosine similarity, do not necessarily correlate with auditory perceptual similarity, which is our primary concern.
-While perceptual measurements exist for speech, these metrics are tailored to speech-specific characteristics and do not translate well to general audio.
-Significant differences exist between speech and general audio regarding frequency bands, harmonic content, and spectral characteristics.
-Most notably, speech typically occupies a narrower frequency range, whereas general audio spans the full audible spectrum and includes diverse timbral qualities, complex textures, and non-harmonic elements that speech-based perceptual metrics cannot adequately capture.
-To the best of our knowledge, established metrics for measuring perceptual similarity in general audio are lacking.
-Therefore, we employ spectrograms to provide intuitive visual analysis in lieu of numerical assessment.
-[^1] All spectrograms presented in this paper are generated using Sonic Visualiser software.
-The corresponding audio is available online.
-[^2] https://github.com/Jng-L/SoundFusion
-(a) Employ inverted latents from our method.
-(b) Employ latents with noise added from forward diffusion.
-(c) Employ noisy latents with stochastic terms preserved from our method.
-Figure 2: Second-order SDE reconstruction with null prompt guidance. Columns 1-3: intervention at t=0, 80, 140. Column 4: original audio through VAE.
+Perceptual similarity is our primary concern for sound fusion.
+For general audio, especially abstract audio defined in our work, there is currently a lack of valid objective perceptual measurements due to unclear semantics, spectral complexity, and psychoacoustic influences.
+There exist perceptual metrics for speech JND ; however, speech occupies a narrow frequency range, whereas general audio spans the full spectrum with diverse timbral qualities, complex textures, and non-harmonic elements that speech metrics cannot adequately capture.
+Therefore, we employ spectrograms for intuitive visual analysis, presenting fusion results and subjective evaluation to assess fusion effects, while we assess signal quality for reconstruction.
 4.1 Original Audio Reconstruction
-To edit a real sample, text prompts serve as critical conditions to guide the sampling process, requiring a balance between identity preservation in latents and responsiveness to prompt guidance.
-Before fusion, we first examine the resistance of latent codes produced by our inversion models to prompt conditioning.
-We use null prompt embeddings in all experiments in this subsection.
+We first examine the capability of our inversion models for reconstructing original audio.
 4.1.1 Necessity of Omitting Stochastic Terms
-Stochastic terms are added back during sampling because they help preserve the potential for editing. Let $\tilde{x}_{t}$ be given by Eq. 6; then the comparisons are as follows:
+Stochastic terms are added back during sampling because they help preserve the potential for editing.
+Let $\tilde{x}_{t}$ be given by Eq.
+3.
+We evaluate three different formulations for $x_{t}$: (a) our proposed inversion model, (b) an approximation of prior work edit-friendly adapted to our case, and (c) preserving the stochastic term in our inversion model.
   1. (a)
@@
     $x_{t}=\tilde{x}_{t}-\sigma_{t}\sqrt{e^{2(\lambda_{t-1}-\lambda_{t})}-1}\epsilon_{t}$
-The first observation is illustrated in Figure 2: any form of noise added during inversion, including retaining stochastic terms or adding noise as in forward diffusion, results in a very noisy sampling phase.
-This observation demonstrates that omitting noise in inversion is reasonable, and that approximating inversion with diffusion produces samples of unacceptable quality with DPMSolver++.
-The original audio is reconstructed through VAE encoding and decoding to eliminate the influence of the VAE itself.
-4.1.2 Potential of Initial Latent
+The results illustrate that our approach achieves the best performance among the three methods, where the dashed lines represent the mean performance for each approach in Figure 1.
+Any form of noise added during inversion—whether adding noise as in forward diffusion or retaining stochastic terms—results in a very noisy sampling phase.
+Notably, audio clips were reconstructed through Variational AutoEncoder (VAE) encoding and decoding to eliminate the influence of the VAE itself.
+Figure 1: Audio Quality Metrics Comparison: Three Inversion Approaches.
+4.1.2 Comparisons with Audio-to-Audio Approaches
+We compare our method with existing approaches, including:
+  • Sampling from scratch.
+  • Image-to-image methods from the visual domain, adapted to the Stable Audio Open workflow for audio-to-audio editing.
+  • The audio-to-audio capabilities of Stable Audio 2.0.
+For the open-source audio-to-audio method, we encoded the original audio through a VAE and incorporated it into randomly initialized latent representations.
+To ensure a fair comparison, we evaluated the SDE inversion model independently on the initial latent that deliberately avoided introducing additional adaptive editing schemes during the sampling process.
+We employed common knowledge prompts—those well-represented in pre-training data—as text inputs for all comparative methods.
+Stable Audio 2 is a commercial versionstable_audio_2 whose audio-to-audio function and sampling process remain black boxes.
+Thus, we included it as an additional reference for comparison.
+There are subtle differences among the generated audio from the open-source model, and it is practically impossible to distinguish the original audio components across all four samples (see Figure 2).
+As expected, audio generated by other methods should exhibit greater similarity to the original audio than samples generated from scratch; however, all outputs of Stable Audio Open sounded similar regardless of the original audio content.
 Figure 2: Comparison of audio generated from initial latents under different methods.
-Prompt: ”An airplane takes off.” (a) Original audio through VAE.
+Prompt: "An airplane takes off." (a) Original audio through VAE.
 (b) Sample from scratch.
@@
 (e) Built-in audio-to-audio of Stable Audio 2.0.
-We evaluate the SDE inversion model independently on the initial latent.
-This evaluation is based on the assumption that, apart from the initial latent representation, the model, sampling process, parameters, and other conditions remain the same.
-The accuracy with which the initial latent representation captures the characteristics of the original audio determines whether the interpolations in the generation process can precisely produce sound containing the original audio’s features and the degree of fidelity.
-We compare our method with existing approaches.
-The comparison methods include:
-  • Sampling from scratch.
-  • Image-to-image methods from the visual domain, adapted to the Stable Audio Open workflow for audio-to-audio editing.
-  • The audio-to-audio capabilities of Stable Audio 2.0.
-In the open-source audio-to-audio method, we encode the original audio through a VAE and incorporate it into randomly initialized latent representations.
-To ensure fair comparison, we maintain consistency across all methods using Stable Audio Open, including audio-to-audio, generating from scratch, and our inversion model.
-Since our objective is to examine the potential of initial latent representations, we deliberately avoid introducing additional adaptive editing schemes during the sampling process.
-Specifically, we do not intervene in the inverted latent representation throughout sampling but maintain methodological consistency by solely modifying the initial latent assignment.
-For evaluation, we employ common knowledge prompts, those with sufficient training data during pre-training, as text inputs.
-We apply identical text prompts and original audio (except for generating from scratch) across all comparative methods, ensuring that the same sampler and parameters are applied to maintain experimental validity.
-Stable Audio 2 is a commercial version, and both its built-in audio-to-audio function and the sampling process remain black boxes to us.
-While it claims to have incorporated this functionality on top of the previous version (Evans et al., 2024a), it is unclear whether it employs the same methodology as the open-source approach.
-Additionally, this version demonstrates improved model capabilities compared to the previous version, so we include it as an additional reference point.
-There are subtle differences among the generated audio from the open-source model, and it is practically impossible to distinguish the original sound components across all five samples (see Figure 3).
-As expected, audio generated by other methods should exhibit greater similarity to the original sound than samples generated from scratch; however, all outputs of Stable Audio Open sound similar regardless of the original audio content.
-Though the sounds generated by the open-source and commercial versions vary, both adhere to the text prompt ”An airplane takes off.”
-As discussed in the methodology section, this phenomenon arises due to the noise control coefficient and signal strength coefficient, which make the noise guided by text prompts dominate generation during the early sampling stages.
+This phenomenon arises due to the noise control coefficient and signal strength coefficient, which make the noise guided by text prompts dominate generation during the early sampling stages.
 This is particularly evident in generation from scratch, where the influence of the initial latent representation is relatively weak.
-These findings indicate that additional controls over the sampling trajectory need to be integrated into the sampling process to achieve ideal reconstruction results.
-4.1.3 Latent Intervention Across Sampler Orders
-Figure A.2 demonstrates reconstructed samples when intervening with inverted latent codes at timesteps 0, 80, 120, 140, and 180, compared to the sample $x_{0}$ in the last column.
-The three rows employ first-, second-, and third-order samplers respectively.
-First-order sampling is equivalent to DDIM.
-It is evident that features illustrated in the third-order reconstruction are relatively vague compared to the other two.
-Although we insert the latent at certain timesteps, higher-order methods compute with interpolations, therefore, the influence of these interpolations persists.
-Consequently, the intervened latents do not dominate the reconstruction to the same extent as in first- and second-order sampling.
-At the end of sampling, sigma, the noise scheduler, becomes very small, and the predicted noise cannot compete with the inserted latents that preserve most of the features.
-This indicates that initially, the prompt controls the generation process, where sigma is fairly large.
+These observations indicate that additional controls over the sampling trajectory need to be integrated into the sampling process to achieve ideal fusion results.
 4.2 Reference Audio Reconstruction
-The audio reconstructed through model fine-tuning restores the reference sound fairly completely, but some loss and confusion of details still exist.
-As shown in the spectrograms of Figure 4, the reconstruction process generates a layer of blurry background noise, with the sound’s stylistic features enhanced over this background noise.
-This phenomenon is most evident in the low-frequency portions.
-As we analyzed earlier, the VAE is designed to prioritize audibility and tends to smooth the regions around sampling points, leading to such artifacts in the sound.
-While the sound lacks clarity in details, it is acceptable overall.
-Therefore, we believe that the noise prediction module has acquired the capability to reconstruct text descriptions into audio with reference sound characteristics.
-Figure 4: Spectrograms of (a) reference audio and (b) reconstructed reference audio.
-4.3 Audio Fusion
-We observe that inverted latent representation intervention at timesteps within a certain interval produces audio with optimal fusion results, indicating the existence of a fusion window.
+We fine-tuned Stable Audio Open for reference audio reconstruction.
+While the reconstructed audio restores the reference audio fairly completely, some detail loss and artifacts still persist.
+As shown in the spectrograms in Figure 3, the reconstruction process generates a layer of blurry background noise, with the audio’s stylistic features enhanced against this background noise.
+This phenomenon is most evident in the low-frequency regions.
+The reason for this can be found in Appendix E.
+While the audio lacks detail clarity, it is acceptable overall.
+Therefore, we believe that the denoising module has acquired the capability to reconstruct text descriptions into audio with reference audio characteristics.
+Figure 3: Spectrograms of reference audio (left) and reconstructed reference audio (right).
+4.3 Sound Fusion
+We observed that inverted latent representation intervention at timesteps within a certain interval produced audio with optimal fusion results, indicating the existence of a fusion window.
 Within this window, as the intervention timestep decreases, the fused audio gradually transitions from resembling the reference audio to resembling the original audio in terms of auditory characteristics.
-Specifically, fusion represents a process of mutual balance between the reference sound and the original sound, where identifying the equilibrium point between the two within the window becomes crucial.
-In this section, we present the results of audio fusion and subjective evaluation of the fusion effectiveness.
-(a) Set 1.
-(b) Set 2.
-(c) Set 3.
-(d) Set 4.
-(e) Set 5.
-Figure 6: Sound fusion results. The first, second, and third columns represent the original sound, reference sound, and fused sound, respectively.
-4.3.1 Results
-We used text descriptions as control conditions for the noise prediction network, intervening with inverted latents during the sampling process at regular intervals of 10 timesteps, starting from $t=0$.
-When intervention occurs near the end of sampling, the sound is almost indistinguishable from the original sound, and the reference sound components are difficult to recognize.
-We therefore intervened with latent representations in the middle-to-late stages of the sampling process to achieve optimal fusion results.
-Figure 6 shows fusion examples.
-4.3.2 Perceptual Evaluation of Fusion Effects
+Specifically, fusion represents a process of mutual balance between the reference audio and the original audio, where identifying the equilibrium point between the two within the window becomes crucial.
 We received a total of 48 valid questionnaires, with most respondents frequently participating in music activities.
-Professional and vocational practitioners constitute the largest proportion, as shown in Figure 7.
-Figure 7: Distribution of respondents’ level of music engagement.
-We observed in our experiments that as the intervention timestep decreased, the auditory similarity of the fused sound gradually transitioned from resembling the reference sound to being more similar to the original sound.
-We selected fusion examples where the intervention timesteps are positioned within the fusion window, and the similarity distribution from subjective evaluation aligns well with our experimental observations.
-We used the fusion examples from the previous section to query respondents.
-In Figure 8, ”Audio 1” represents the original audio, ”Audio 2” represents the reference audio, and ”Audio 3” represents the fused audio; the first through fifth audio sets correspond to Set 1 through Set 5 in Figure 6, respectively.
+Music students make up the largest group with 22 participants (45.83%); music professionals and enthusiasts each represent equal portions of the sample, with 12 participants each (25.0%); only 4.17% are infrequent listeners.
+We selected fusion examples where the intervention timesteps are positioned within the fusion interval.
+In Figure 4, "Audio 1" represents the original audio, "Audio 2" represents the reference audio, and "Audio 3" represents the fused audio; the first through fifth audio sets correspond to Set 1 through Set 5 in Figure D.2, respectively.
 Overall, respondents rated the fusion effects favorably, indicating that we successfully achieved the goal of sound fusion.
@@
 In this paper, we propose inversion models based on DPMSolver++ with two variations: SDE and ODE.
-We make the integral term as constant, thereby avoiding the circular dependency of expanding the noise prediction within the sampling formula.
-Additionally, the deterministic formulation allows for straightforward computation of the noise map at each inversion step.
-Our approach retains the flexibility of prompt guidance for both positive and negative guidance.
-It is computationally efficient.
+We treat the integral term as a constant, thereby avoiding the circular dependency (detailed in Appendix A) that arises from expanding the noise prediction within the sampling formula.
+Additionally, the deterministic formulation enables straightforward computation of the noise map at each inversion step.
 We explain the necessity of removing stochastic terms from the inversion model, demonstrating that approximating inversion with forward diffusion is not applicable for DPMSolver++ sampling phase.
-Previous inversion approaches typically computed only the initial latent, and our experiments show that the initial latent alone is insufficient to capture features of the original sample.
-Therefore, for editing applications, it must be combined with additional control approaches, such as optimized prompt embeddings that control the sampling trajectory, or by intervening in the sampling phase as we propose.
-In the latter case, we do not necessarily need to invert all the way to the final timestep.
-Inversion is employed to reconstruct the auditory features of the original sound, while fine-tuning helps preserve the features of the reference sound through prompting.
+Our experiments show that the initial latent alone is insufficient to preserve the features of inverted audio against text guidance in sampling.
+Therefore, it must be combined with additional control mechanisms, such as optimized prompt embeddings that guide the sampling trajectory, or intervention in the sampling phase as we propose for fusion purposes.
+In the latter approach, we do not necessarily need to invert all the way to the final timestep.
+Inversion is employed to reconstruct the auditory features of the original audio, while fine-tuning helps preserve the features of the reference audio through prompting.
 An optimal interval exists during which intervention achieves effective fusion.
-As the timestep decreases, the auditory features of the fused sound shift from those of the reference sound to those of the original sound.
+As the timestep decreases, the auditory features of the fused audio shift from those of the reference audio to those of the original audio.
 Higher-order samplers provide a wider fusion interval.
-Due to the lack of perceptual metrics for general audio, and because our sound fusion approach does not conform to specific physical constraints, as indicated by participant feedback in Appendix B, objective evaluation of our fusion results presents significant challenges.
+Due to the lack of perceptual metrics for general audio, and because our audio fusion approach does not conform to specific physical constraints, as indicated by participant feedback in Appendix D, objective evaluation of our fusion results presents significant challenges.
 Therefore, we primarily rely on subjective auditory evaluation.
 The results validate the effectiveness of the sound fusion and inversion methods proposed in this work.
-Appendix A Latent Intervention across Sampler Orders: Additional Results
-This section provides comprehensive visual results for the latent intervention experiments across different sampler orders discussed in Section 4.1.3.
-Figure A.2 shows detailed spectrograms comparing first-, second-, and third-order sampling with interventions at various timesteps.
+Appendix A Predicted Noise Introduces Circular Dependency
+Lu et al.
+DPMSolver++ proved that many high-order samplers are unstable at large guided scales, which amplified the derivatives of the noise prediction model and affect the convergence range of the ODE solver.
+The authors proposed replacing the noise prediction term with the model output $x_{\theta}$ to achieve more stable convergence.
+Previous research performed inversion on predicted noise.
+However, expanding the model output with the predicted noise expression results in circular dependency.
+To address this circular dependency issue intuitively, we substitute the general formula in Eq.
+2 into the second-order SDE case and target the latent at the next timestep $r$ as below:
+$\boldsymbol{x}_{r}=\frac{2r_{1}(\sigma_{r}^{2}+1)}{\alpha_{t}(1-e^{-2h})}\left(\boldsymbol{x}_{t}-\frac{\sigma_{t}}{\sigma_{s}}e^{-h}\boldsymbol{x}_{s}+\frac{\alpha_{t}(1-2r_{1})(1-e^{-2h})}{2r_{1}}x_{\theta}^{(s)}+\frac{\sigma_{r}\alpha_{t}(1-e^{-2h})}{2r_{1}\sqrt{\sigma_{r}^{2}+1}}\epsilon^{(r)}_{\theta}\right)$
+In this example, $\epsilon_{\theta}$ takes $\boldsymbol{x}_{r}$ as its input, which is the variable we are solving for on the left side.
+This circular dependency occurs in first- and higher-order methods for both SDE and ODE samplers.
+Fine-tuning on latent $\boldsymbol{x}_{r}$ becomes difficult because the outer scaling factor grows extremely large during the latter stages of inversion when the maximum $\sigma$ is set to its default value.
+Consequently, small error accumulation leads to numerically unstable training.
+Appendix B Derived Noise Maps
+We have inversion models for both SDE and ODE variants of DPMSolver++. We can further derive the noise map at $t$ from Eq. 1:
+$\boldsymbol{\epsilon}^{(t)}=\frac{\boldsymbol{x}_{t}-\alpha_{t}\boldsymbol{x}_{0}}{\sigma_{t}}$
+We demonstrate the relationships between inverted latent interpolations and derived noise maps, as well as the interpolations in the sampling process that utilize the inverted latent and its corresponding noise map at each step in Figure B.2.
+(a) Interpolations during inversion (time steps increasing from left to right).
+(b) Noise maps computed from latent interpolations at the same timestep of the inversion.
+(c) Restoration of the original audio through sampling using inverted latent interpolations and noise maps (timesteps decreasing from left to right).
+Figure B.2: Sampling from inverted latents and derived noise maps, such that the sampling trajectory coincides with the inversion trajectory.
+Appendix C Latent Intervention Across Sampler Orders
+Figure C.2 demonstrates reconstructed samples when intervening with inverted latent codes at timesteps 0, 80, 120, 140, and 180, compared to the sample $\boldsymbol{x}_{0}$ in the last column.
+The three rows employ first-, second-, and third-order samplers respectively.
+First-order sampling is equivalent to DDIM.
+It is evident that features illustrated in the third-order reconstruction are relatively vague compared to the other two.
+Although we insert the latent at certain timesteps, higher-order methods compute with interpolations, therefore, the influence of these interpolations persists.
+Consequently, the intervened latents do not dominate the reconstruction to the same extent as in first- and second-order sampling.
+At the end of sampling, sigma, the noise scheduler, becomes very small, and the predicted noise cannot compete with the inserted latents that preserve most of the features.
+This indicates that initially, the prompt controls the generation process, where sigma is fairly large.
 (a) Audio reconstructed by first-order sampling.
@@
 (c) Audio reconstructed by third-order sampling.
-Figure A.2: SDE reconstruction after inversion with null prompt guidance. Columns 1-5: intervention at t=0, 80, 120, 140, 160. Column 6: original audio through VAE.
-Appendix B Subjective Analysis on Fusion Results
+Figure C.2: SDE reconstruction after inversion with null prompt guidance. Columns 1-5: intervention at t=0, 80, 120, 140, 160. Column 6: original audio through VAE.
+Appendix D Subjective Analysis on Fusion Results
 Fusion is achieved through music generation models, making it difficult to control granularity at the physical or acoustic level.
-Professional electronic music analysis reveals distinct fusion patterns across the examples.
-In the first example, the original audio concentrates in high frequencies, while the reference audio shows evenly distributed frequencies with prominent highs.
-The fused audio resembles the reference audio’s frequency distribution but adopts the ”melodic direction” of the original audio.
+This section presents professional electronic music analysis that reveals distinct fusion patterns across the examples.
+In the first example of Figure D.2, the original audio concentrates in high frequencies, while the reference audio shows evenly distributed frequencies with prominent highs.
+The fused audio resembles the reference audio’s frequency distribution but retains the "melodic direction" of the original audio.
 In the second example, the original audio exhibits evenly distributed frequencies with prominent highs and minimal reverb; the reference audio has obvious melodic characteristics with concentrated high tones and some reverb; the fused audio has significant reverb and combines melodies from both sources.
@@
 In the fifth example, the original audio is primarily rhythmic with percussion with layered melody; the reference audio is pure noise with spatial characteristics; the fused audio combines the rhythm and melody of the original with the spatial characteristics of the reference, amplifying these effects.
-Appendix C Derived Noise Maps
-At the end of Section 3.3, we derived the noise map at timestep $t$ given $x_{0}$ and $x_{t}$ in Eq.
-9.
-Below, we demonstrate the relationships between inverted latent interpolations and derived noise maps, as well as the interpolations in the sampling process that utilize the inverted latent and its corresponding noise map at each step:
-(a) Interpolations during inversion (time steps increasing from left to right).
-(b) Noise maps computed from latent interpolations at the same timestep of the inversion.
-(c) Restoration of the original sound through sampling using inverted latent interpolations and noise maps (timesteps decreasing from left to right).
-Figure C.2: Sampling from inverted latents and derived noise maps, such that the sampling trajectory coincides with the inversion trajectory.
-Appendix D VAE Blurs Acoustic Details
-Initially, the audio is encoded by the Variational AutoEncoder (VAE) to derive $x_{0}$.
-Figure D.1 presents a comparative analysis: the left panel depicts the spectrogram of the original audio signal; the center shows the spectrogram of audio reconstructed through VAE encoding-decoding processes; the right shows audio reconstructed from inverted $x_{t}$ using null prompt guidance.
+(a) Set 1.
+(b) Set 2.
+(c) Set 3.
+(d) Set 4.
+(e) Set 5.
+Figure D.2: Audio fusion results. The first, second, and third columns represent the original audio, reference audio, and fused audio, respectively.
+Appendix E VAE Blurs Acoustic Details
+Initially, the audio is encoded by the VAE to derive $x_{0}$.
+Figure E.1 presents a comparative analysis.
 While GANs typically produce high-definition samples, VAEs generate comparatively smoother outputs.
@@
 Consequently, evaluations in this paper compare reconstructions through VAE encoding-decoding with original samples, rather than comparing with the original samples directly.
-Figure D.1: Spectrogram of the original audio (left), spectrogram of audio reconstructed through VAE encoding-decoding (center), and spectrogram of audio reconstructed by sampling from the inverted latent intervened at step 100/ 200 (right).
+Figure E.1: Spectrogram of the original audio (left), spectrogram of audio reconstructed through VAE encoding-decoding (center), and spectrogram of audio reconstructed from inverted $x_{t}$ at step 100/ 200 using null prompt guidance (right).